# Adapting LLaMA Decoder to Vision Transformer

### v1

@@ Abstract
This work examines whether decoder-only Transformers such as LLaMA, which were originally designed for large language models (LLMs), can be adapted to the computer vision field. We first “LLaMAfy" a standard ViT step-by-step to align with LLaMA’s architecture, and find that directly applying a casual mask to the self-attention brings an attention collapse issue, resulting in the failure to the network training. We suggest to reposition the class token behind the image tokens with a post-sequence class token technique to overcome this challenge, enabling causal self-attention to efficiently capture the entire image’s information. Additionally, we develop a soft mask strategy that gradually introduces a casual mask to the self-attention at the onset of training to facilitate the optimization behavior. The tailored model, dubbed as image LLaMA (iLLaMA), is akin to LLaMA in architecture and enables direct supervised learning. Its causal self-attention boosts computational efficiency and learns complex representation by elevating attention map ranks. iLLaMA rivals the performance with its encoder-only counterparts, achieving 75.1% ImageNet top-1 accuracy with only 5.7M parameters. Scaling the model to $\sim$310M and pre-training on ImageNet-21K further enhances the accuracy to 86.0%. Extensive experiments demonstrate iLLaMA’s reliable properties: calibration, shape-texture bias, quantization compatibility, ADE20K segmentation and CIFAR transfer learning. We hope our study can kindle fresh views to visual model design in the wave of LLMs. Pre-trained models and codes are available here.
@@ 1 Introduction
The year 2024 saw the meteoric rise of large language models (LLMs), as well as the 4th anniversary of the Vision Transformer (ViT) dosovitskiy2020image . Born in 2020, ViT has revolutionized the computer vision field by adapting the Transformer vaswani2017attention , which was originally designed for natural language, to process images as non-overlapping patches. Notably, ViT design was influenced by the prevailing encoder-only text Transformers at the time, such as BERT devlin2018bert , etc. Accordingly, ViT is allowed to borrow the encoder-only design, i.e., self-attention do not use any casual mask. As a result, advanced vision backbones touvron2021training ; wang2021pyramid and training paradigms bao2021beit ; he2022masked have since followed such convention by default.
@@
In the era of LLMs, could decoder-only Transformers be revitalized in the vision field? In this study, we demonstrate that through straightforward supervised learning, LLaMA architecture itself can process input images with simple yet crucial modifications. We start by modifying a standard encoder-only ViT (e.g., ViT-T/16), progressively adapting its components to align with those in LLaMA. In practice, we observe an attention collapse issue, i.e., the training loss fails to converge by directly adding a casual mask to the attention map. The causal mask restricts the class token from accessing the image’s global information, thereby hindering the optimization of the training loss. To this end, we propose a post-sequence class token technique, repositioning the class token to the end of image tokens (details in Section 3.3). As a result, casual mask can keep the attention score between the class token and others, allowing the model to optimize stably. We also evaluate the advantages of the casual self-attention in reducing computational complexity and enhancing the attention map rank.
@@
Figure 1: Left: iLLaMA architecture. Right: the design roadmap. Colored and gray bars represent the results of the tiny and base regimes, with the red line depicting the training loss of the tiny regime. iLLaMA strives to process visual tokens using standard LLaMa components, e.g., casual-self-attention. The proposed PS [cls] and soft mask strategy help overcome training challenges. Block details of ViT dosovitskiy2020image , VisionLLaMA chu2024visionllama , and our iLLaMA is compared in Figure 5 in Appendix A.
@@
Moreover, we explore several training techniques for the proposed casual Transformer. When observing things, humans start by broadly catching global connections, then narrow down to focus on specifics. Motivated by this, we develop a soft mask approach – bi-directional self-attention degenerates to a casual self-attention at the onset of training – to further boost the network performance. Soft mask does not alter the casual self-attention during inference but improves the initial training behavior of the network. (details in Section 3.6). We illustrate different types of masks in Figure 3.
@@
Equipped with such modifications, we propose a non-autoregressive decoder-only vision Transformer with casual self-attention inside, dubbed image LLaMA (iLLaMA), as shown in Figure 1. We provide a block level comparison of ViT dosovitskiy2020image , VisionLLaMA chu2024visionllama , and our iLLaMA in Figure 5 in Appendix A. iLLaMA process visual input with our slightly modified casual self-attention. We conduct a comprehensive evaluation of iLLaMA’s properties, including ImageNet-1K classification deng2009imagenet , calibration, shape-texture bias, quantization compatibility, ADE20K semantic segmentation zhou2019semantic , and CIFAR transfer learning krizhevsky2009learning . Experimental results show that iLLaMA delivers favorable and reliable performance to its encoder-only counterparts (i.e., ViT, VisionLLaMA), while maintaining a pure decoder design. More importantly, a spectral analysis on the attention map empirically shows that compared to bi-directional counterparts, casual self-attention has a higher rank (see Figure 4), which allows for learning complex image representation. We hope our work to inspire a re-evaluation of vision backbone design in the era of LLMs and provide fresh insights for their architectural unification.
@@ 2 Preliminaries and Motivation
where $i,j\in[1,N]$, ${\bf A}\in\mathbb{R}^{N\times N}$, ${\bf M}\in\mathbb{R}^{N\times N}$ denote the attention map and mask. ${\bf P}\in\mathbb{R}^{N\times N}$, ${\bf Q}\in\mathbb{R}^{N\times N}$ are masks in the encoder and decoder, respectively. For a casual self-attention, we have ${\bf M}={\bf Q}$. Such design allows subsequent tokens only attend to the preceding ones, but not vice versa. For a bi-directional self-attention, we have ${\bf M}={\bf P}$, ensuring mutual visibility for each token.
@@ 3.2 Normalization Layer
   (a) casual mask         (b) casual mask w/ PS [cls]        (c) modified casual mask
Figure 2: Illustration of different masks in self-attention. From left to right: mask in casual self-attention, mask in casual self-attention with the PS [cls] method, and a modified casual mask.
3.3 Casual Self-Attention Leads to Attention Collapse
Attention collapse issue. As a common practice for Transformer decoders, the key component for casual self-attention is the casual mask, i.e., a lower triangular mask matrix, illustrated in Eq. 1 and Figure 2(a). With such, each token can get the attention score of all its previous ones. We add the casual mask to our network via a non-autoregressive way. The reason is that visual perception tasks, unlike text generation, require only inference once. As a result, we observe that the training loss fails to converge in both ViT-T/16 and ViT-B/16 regimes (line 1 in Table 1). We posit that such issue stems from the influence of the lower triangular matrix, which prevents the class token from “seeing" other image tokens. As illustrated in Figure 2(a), when the class token is positioned at the start of the patch embedding, its attention score for all other image tokens gets zero due to a casual mask. We term this occurrence as the attention collapse issue, which leads to a loss of connection between the class token and other image patches, thereby hindering the optimization of the network.
Table 1: Results of PS [cls] and the modified casual mask. Training converges in both settings.
@@ 3.3 Casual Self-Attention Leads to Attention Collapse
Post-sequence class token (PS [cls]). The attention collapse issue stems from the inappropriate placement of the token. To this end, we suggest a PS [cls] strategy, by placing it at the end of the token sequence, without changing the casual mask, as shown in Figure 2(b) and Figure 1. Such modification ensures that the class token can achieve global information about all image tokens, while maintaining a casual self-attention property. As a result, we observe that the attention collapse issue is eliminated and the training process starts to stabilize, leading the network performance to $71.9\%$ for ViT-T/16 and $80.6\%$ for ViT-B/16 regime, respectively (line 2 in Table 1).
To test our hypothesis about the reason of the attention collapse issue, we also explore a mask setting in Figure 2(c). In this setting, we do not change the position of the class token. Instead, we unmask the first row of the mask (i.e., attention score of the class token) on the basis of the casual self-attention, termed as “modified casual mask". Ablation results (line 3 in Table 1) shows that both settings can solve the attention collapse issue as expected, and the “modified casual mask" leads to a better $72.5\%$ accuracy for ViT-T/16 regime, validating our hypothesis about the reason. Although the results do not surpass the performance of bi-directional counterpart, they demonstrate the potential for optimizing casual self-attention in a decoder-only image model. We also observe that the PS [cls] method yields higher accuracy with a slightly larger training loss for ViT-B/16 regime, suggesting lower overfitting.
We will employ casual self-attention with the proposed PS [cls] method in each block.
@@ 3.5 Data Augmentation
Mixup zhang2017mixup and cutmix yun2019cutmix that we used to train our iLLaMA (0.8 and 1.0), are borrowed from DeiT touvron2021training ’s recipe. Unlike the bi-directional self-attention used in DeiT, casual self-attention affects the connection between image tokens. Meanwhile, these two hyper-parameters affect the content of the input image, which further influences the subsequent embedding. Thus, we reevaluate their impact on iLLaMA optimization. Specifically, we discover that a combination of 0.1 mixup and 0.1 cutmix improves the performance of the iLLaMA-T/16 to $74.3\%$, whereas a combination of 0.95 and 1.0 leads the iLLaMA-B/16 to a $81.3\%$ accuracy. Other ablations are detailed in Section 4.1.
@@ 3.6 Soft Mask Strategy: Optimization Improvement
When observing objects, humans tend to perceive broad connections, then focus on specific details. Motivated by this, we propose a soft mask technique to improve the model’s optimization: starting with bi-directional self-attentions in the early training epochs and gradually shifting completely to casual self-attentions as the optimization goes. Specifically, self-attention can be formulated as:
@@
Figure 3: Left: soft mask gradually transitions from a bi-directional mask into a casual mask during training through a constant or linear schedule. Right: ablation training loss and test accuracy.
where $i,j\in[1,N]$, ${\bf S}\in\mathbb{R}^{N\times N}$ denotes the soft mask, which is defined as a linear combination of a bi-directional mask ${\bf B}$ and a casual mask ${\bf C}$. $\alpha$ is the hyper-parameter controlling the mask configuration, i.e., soft mask degenerates into ${\bf B}$ or ${\bf C}$ when $\alpha=1$ or $\alpha=0$, respectively. As illustrated in Figure 3(a), $\alpha$ involves three related hyper-parameters: 1) scheme: how $\alpha$ drops from 1 to 0: we try a linear or a constant scheme. 2) cutoff epochs: when will $\alpha$ drops to 0. 3) learning rate (lr) warmup he2016deep ; goyal2017accurate : this hyper-parameter overlaps with the duration of soft mask. We initially set the lr warmup epochs at 50, consistent with previous settings. When using a linear scheme with 50 and 25 cutoff epochs, we observe an improvement in performance for both iLLaMA-T/16 and iLLaMA-B/16 models, achieving $74.9\%$ and $81.6\%$, respectively. Ablations are detailed in Section. 4.1. We plot the training curve of the iLLaMA-T/16 in Figure 3(b), using a constant scheme with 50 cutoff epochs. When soft mask ends, we observe that although there was a sharp drop in accuracy, the model ends up achieving better performance. Similar case of the iLLaMA-B/16 are shown in Appendix E. Additionally, we discover that a lower learning rate warmup helps iLLaMA-T/16 achieve $75.0\%$ accuracy, by using a constant scheme with 50 cutoff epochs. However, such approach leads larger iLLaMA-B/16 training to fail to converge. As a result, we only used the lower learning rate warmup for iLLaMA-T/16. Notably, the final training loss within both iLLaMA-T/16 and iLLaMA-B/16 decreases when using soft masks, suggesting a mitigation of the potential underfitting concern.
3.7 Analysis of Casual Self-attention
Next, we analyze the advantages of using casual self-attention in iLLaMA, in terms of computational efficiency and expressive ability of visual representation through the lens of attention map rank.
Computational Complexity. We compare the efficiency of casual self-attention and bi-directional baseline. For a self-attention with a sequence length $N$ and embedding dimension $D$, FLOPs are reported in Table 2 (RoPE is not involved as only attention computations are considered). Casual self-attention, in light of the lower triangular property of its attention map, slightly reduces the FLOPs compared to the bi-directional baseline — the degree of reduction grows as sequence length increases.
Table 2: Computational complexity results. Casual mask slightly reduces FLOPs required in the self-attention.
Type Bi-directional Casual
@@ 3.7 Analysis of Casual Self-attention
Attention map rank. We examine the representation learning power of casual attention through a spectrum analysis. Following wang2020linformer ; shu2021adder , we perform singular value decomposition on the attention maps of the pre-trained ViT-T/16 and iLLaMA-T/16 models. Next, we sort the singular values and plot a curve illustrating the relationship between the cumulative normalized singular values and matrix indices. The results are conducted using $30$ images randomly selected from the ImageNet-1K validation set. As shown in Figure 4, the curve of ViT exhibits concave function characteristics, while the curve of iLLaMA is close to a linear function, indicating a more uniform distribution of singular values in iLLaMA’s attention map. Approximating the matrix rank by the index at which the cumulative normalized singular value reaches 0.8, we observe that the index value of iLLaMA is about 48 higher than that of ViT ($\sim$129-th v.s. $\sim$81-th). Under such premise, compared to ViT, the attention map of iLLaMA can be approximated with a certain error by a higher-rank matrix. Accordingly, the rank of the attention map may affect the expressive capabilities of the learned representations dong2021attention , suggesting that the casual self-attention in iLLaMA has the potential to learn complex visual representations, as empirically demonstrated in Section 4.2. Detailed results for different layers and heads are provided in Appendix D.
@@
Closing remarks. So far, we have finished the design roadmap of iLLaMA through architectural and training strategy modification. iLLaMA, a decoder-only Transformer, shows advantages in computational complexity and attention map rank through its casual self-attention mechanism. Notably, while all components of iLLaMA are essentially derived from LLaMA, relying only on them is insufficient for an effective weight optimization, as demonstrated in Section 3.3. In fact, the proposed PS [cls] and soft mask strategy effectively address this issue and assist in iLLaMA training. However, to achieve a comprehensive understanding of iLLaMA’s properties, some useful evaluation should be conducted: 1) Scalability for large model capacities (>300M parameters) and dataset sizes (>10M training images, e.g., ImageNet-21K). 2) Other practical evaluation dimensions, such as model calibration, shape-texture bias, downstream task performance, quantization compatibility, discussed below.

### v2

@@ Abstract
This work examines whether decoder-only Transformers such as LLaMA, which were originally designed for large language models (LLMs), can be adapted to the computer vision field. We first “LLaMAfy" a standard ViT step-by-step to align with LLaMA’s architecture, and find that directly applying a causal mask to the self-attention brings an attention collapse issue, resulting in the failure to the network training. We suggest to reposition the class token behind the image tokens with a post-sequence class token technique to overcome this challenge, enabling causal self-attention to efficiently capture the entire image’s information. Additionally, we develop a soft mask strategy that gradually introduces a causal mask to the self-attention at the onset of training to facilitate the optimization behavior. The tailored model, dubbed as image LLaMA (iLLaMA), is akin to LLaMA in architecture and enables direct supervised learning. Its causal self-attention boosts computational efficiency and learns complex representation by elevating attention map ranks. iLLaMA rivals the performance with its encoder-only counterparts, achieving 75.1% ImageNet top-1 accuracy with only 5.7M parameters. Scaling the model to $\sim$310M and pre-training on ImageNet-21K further enhances the accuracy to 86.0%. Extensive experiments demonstrate iLLaMA’s reliable properties: calibration, shape-texture bias, quantization compatibility, ADE20K segmentation and CIFAR transfer learning. We hope our study can kindle fresh views to visual model design in the wave of LLMs. Pre-trained models and codes are available here.
@@ 1 Introduction
The year 2024 saw the meteoric rise of large language models (LLMs), as well as the 4th anniversary of the Vision Transformer (ViT) dosovitskiy2020image . Born in 2020, ViT has revolutionized the computer vision field by adapting the Transformer vaswani2017attention , which was originally designed for natural language, to process images as non-overlapping patches. Notably, ViT design was influenced by the prevailing encoder-only text Transformers at the time, such as BERT devlin2018bert , etc. Accordingly, ViT is allowed to borrow the encoder-only design, i.e., self-attention do not use any causal mask. As a result, advanced vision backbones touvron2021training ; wang2021pyramid ; guo2022cmt and training paradigms bao2021beit ; he2022masked have followed such convention by default.
@@
In the era of LLMs, could decoder-only Transformers be revitalized in the vision field? In this study, we demonstrate that through straightforward supervised learning, LLaMA architecture itself can process input images with simple yet crucial modifications. We start by modifying a standard encoder-only ViT (e.g., ViT-T/16), progressively adapting its components to align with those in LLaMA. In practice, we observe an attention collapse issue, i.e., the training loss fails to converge by directly adding a causal mask to the attention map. The causal mask restricts the class token from accessing the image’s global information, thereby hindering the optimization of the training loss. To this end, we propose a post-sequence class token technique, repositioning the class token to the end of image tokens (details in Section 3.3). As a result, causal mask can keep the attention score between the class token and others, allowing the model to optimize stably. We also evaluate the advantages of the causal self-attention in reducing computational complexity and enhancing the attention map rank.
@@
Figure 1: Left: iLLaMA architecture. Right: the design roadmap. Colored and gray bars represent the results of the tiny and base regimes, with the red line depicting the training loss of the tiny regime. iLLaMA strives to process visual tokens using standard LLaMa components, e.g., causal self-attention. The proposed PS [cls] and soft mask strategy help overcome training challenges. Block details of ViT dosovitskiy2020image , VisionLLaMA chu2024visionllama , and our iLLaMA is compared in Figure 5 in Appendix A.
@@
Moreover, we explore several training techniques for the proposed causal Transformer. When observing things, humans start by broadly catching global connections, then narrow down to focus on specifics. Motivated by this, we develop a soft mask approach – bi-directional self-attention degenerates to a causal self-attention at the onset of training – to further boost the network performance. Soft mask does not alter the causal self-attention during inference but improves the initial training behavior of the network. (details in Section 3.6). We illustrate different types of masks in Figure 3.
@@
Equipped with such modifications, we propose a non-autoregressive decoder-only vision Transformer with causal self-attention inside, dubbed image LLaMA (iLLaMA), as shown in Figure 1. We provide a block level comparison of ViT dosovitskiy2020image , VisionLLaMA chu2024visionllama , and our iLLaMA in Figure 5 in Appendix A. iLLaMA process visual input with our slightly modified causal self-attention. We conduct a comprehensive evaluation of iLLaMA’s properties, including ImageNet-1K classification deng2009imagenet , calibration, shape-texture bias, quantization compatibility, ADE20K semantic segmentation zhou2019semantic , and CIFAR transfer learning krizhevsky2009learning . Experimental results show that iLLaMA delivers favorable and reliable performance to its encoder-only counterparts (i.e., ViT, VisionLLaMA), while maintaining a pure decoder design. More importantly, a spectral analysis on the attention map empirically shows that compared to bi-directional counterparts, causal self-attention has a higher rank (see Figure 4), which allows for learning complex image representation. We hope our work to inspire a re-evaluation of vision backbone design in the era of LLMs and provide fresh insights for their architectural unification.
@@ 2 Preliminaries and Motivation
where $i,j\in[1,N]$, ${\bf A}\in\mathbb{R}^{N\times N}$, ${\bf M}\in\mathbb{R}^{N\times N}$ denote the attention map and mask. ${\bf P}\in\mathbb{R}^{N\times N}$, ${\bf Q}\in\mathbb{R}^{N\times N}$ are masks in the encoder and decoder, respectively. For a causal self-attention, we have ${\bf M}={\bf Q}$. Such design allows subsequent tokens only attend to the preceding ones, but not vice versa. For a bi-directional self-attention, we have ${\bf M}={\bf P}$, ensuring mutual visibility for each token.
@@ 3.2 Normalization Layer
   (a) causal mask         (b) causal mask w/ PS [cls]        (c) modified causal mask
Figure 2: Illustration of different masks in self-attention. From left to right: mask in causal self-attention, mask in causal self-attention with the PS [cls] method, and a modified causal mask.
3.3 causal Self-Attention Leads to Attention Collapse
Attention collapse issue. As a common practice for Transformer decoders, the key component for causal self-attention is the causal mask, i.e., a lower triangular mask matrix, illustrated in Eq. 1 and Figure 2(a). With such, each token can get the attention score of all its previous ones. We add the causal mask to our network via a non-autoregressive way. The reason is that visual perception tasks, unlike text generation, require only inference once. As a result, we observe that the training loss fails to converge in both ViT-T/16 and ViT-B/16 regimes (line 1 in Table 1). We posit that such issue stems from the influence of the lower triangular matrix, which prevents the class token from “seeing" other image tokens. As illustrated in Figure 2(a), when the class token is positioned at the start of the patch embedding, its attention score for all other image tokens gets zero due to a causal mask. We term this occurrence as the attention collapse issue, which leads to a loss of connection between the class token and other image patches, thereby hindering the optimization of the network.
Table 1: Results of PS [cls] and the modified causal mask. Training converges in both settings.
@@
Post-sequence class token (PS [cls]). The attention collapse issue stems from the inappropriate placement of the token. To this end, we suggest a PS [cls] strategy, by placing it at the end of the token sequence, without changing the causal mask, as shown in Figure 2(b) and Figure 1. Such modification ensures that the class token can achieve global information about all image tokens, while maintaining a causal self-attention property. As a result, we observe that the attention collapse issue is eliminated and the training process starts to stabilize, leading the network performance to $71.9\%$ for ViT-T/16 and $80.6\%$ for ViT-B/16 regime, respectively (line 2 in Table 1).
To test our hypothesis about the reason of the attention collapse issue, we also explore a mask setting in Figure 2(c). In this setting, we do not change the position of the class token. Instead, we unmask the first row of the mask (i.e., attention score of the class token) on the basis of the causal self-attention, termed as “modified causal mask". Ablation results (line 3 in Table 1) shows that both settings can solve the attention collapse issue as expected, and the “modified causal mask" leads to a better $72.5\%$ accuracy for ViT-T/16 regime, validating our hypothesis about the reason. Although the results do not surpass the performance of bi-directional counterpart, they demonstrate the potential for optimizing causal self-attention in a decoder-only image model. We also observe that the PS [cls] method yields higher accuracy with a slightly larger training loss for ViT-B/16 regime, suggesting lower overfitting.
We will employ causal self-attention with the proposed PS [cls] method in each block.
@@ 3.5 Data Augmentation
Mixup zhang2017mixup and cutmix yun2019cutmix that we used to train our iLLaMA (0.8 and 1.0), are borrowed from DeiT touvron2021training ’s recipe. Unlike the bi-directional self-attention used in DeiT, causal self-attention affects the connection between image tokens. Meanwhile, these two hyper-parameters affect the content of the input image, which further influences the subsequent embedding. Thus, we reevaluate their impact on iLLaMA optimization. Specifically, we discover that a combination of 0.1 mixup and 0.1 cutmix improves the performance of the iLLaMA-T/16 to $74.3\%$, whereas a combination of 0.95 and 1.0 leads the iLLaMA-B/16 to a $81.3\%$ accuracy. Other ablations are detailed in Section 4.1.
@@ 3.6 Soft Mask Strategy: Optimization Improvement
When observing objects, humans tend to perceive broad connections, then focus on specific details. Motivated by this, we propose a soft mask technique to improve the model’s optimization: starting with bi-directional self-attentions in the early training epochs and gradually shifting completely to causal self-attentions as the optimization goes. Specifically, self-attention can be formulated as:
@@
Figure 3: Left: soft mask gradually transitions from a bi-directional mask into a causal mask during training through a constant or linear schedule. Right: ablation training loss and test accuracy.
where $i,j\in[1,N]$, ${\bf S}\in\mathbb{R}^{N\times N}$ denotes the soft mask, which is defined as a linear combination of a bi-directional mask ${\bf B}$ and a causal mask ${\bf C}$. $\alpha$ is the hyper-parameter controlling the mask configuration, i.e., soft mask degenerates into ${\bf B}$ or ${\bf C}$ when $\alpha=1$ or $\alpha=0$, respectively. As illustrated in Figure 3(a), $\alpha$ involves three related hyper-parameters: 1) scheme: how $\alpha$ drops from 1 to 0: we try a linear or a constant scheme. 2) cutoff epochs: when will $\alpha$ drops to 0. 3) learning rate (lr) warmup he2016deep ; goyal2017accurate : this hyper-parameter overlaps with the duration of soft mask. We initially set the lr warmup epochs at 50, consistent with previous settings. When using a linear scheme with 50 and 25 cutoff epochs, we observe an improvement in performance for both iLLaMA-T/16 and iLLaMA-B/16 models, achieving $74.9\%$ and $81.6\%$, respectively. Ablations are detailed in Section. 4.1. We plot the training curve of the iLLaMA-T/16 in Figure 3(b), using a constant scheme with 50 cutoff epochs. When soft mask ends, we observe that although there was a sharp drop in accuracy, the model ends up achieving better performance. Similar case of the iLLaMA-B/16 are shown in Appendix E. Additionally, we discover that a lower learning rate warmup helps iLLaMA-T/16 achieve $75.0\%$ accuracy, by using a constant scheme with 50 cutoff epochs. However, such approach leads larger iLLaMA-B/16 training to fail to converge. As a result, we only used the lower learning rate warmup for iLLaMA-T/16. Notably, the final training loss within both iLLaMA-T/16 and iLLaMA-B/16 decreases when using soft masks, suggesting a mitigation of the potential underfitting concern.
3.7 Analysis of causal Self-attention
Next, we analyze the advantages of using causal self-attention in iLLaMA, in terms of computational efficiency and expressive ability of visual representation through the lens of attention map rank.
Computational Complexity. We compare the efficiency of causal self-attention and bi-directional baseline. For a self-attention with a sequence length $N$ and embedding dimension $D$, FLOPs are reported in Table 2 (RoPE is not involved as only attention computations are considered). causal self-attention, in light of the lower triangular property of its attention map, slightly reduces the FLOPs compared to the bi-directional baseline — the degree of reduction grows as sequence length increases.
Table 2: Computational complexity results. causal mask slightly reduces FLOPs required in the self-attention.
Type Bi-directional causal
@@ 3.7 Analysis of causal Self-attention
Attention map rank. We examine the representation learning power of causal attention through a spectrum analysis. Following wang2020linformer ; shu2021adder , we perform singular value decomposition on the attention maps of the pre-trained ViT-T/16 and iLLaMA-T/16 models. Next, we sort the singular values and plot a curve illustrating the relationship between the cumulative normalized singular values and matrix indices. The results are conducted using $30$ images randomly selected from the ImageNet-1K validation set. As shown in Figure 4, the curve of ViT exhibits concave function characteristics, while the curve of iLLaMA is close to a linear function, indicating a more uniform distribution of singular values in iLLaMA’s attention map. Approximating the matrix rank by the index at which the cumulative normalized singular value reaches 0.8, we observe that the index value of iLLaMA is about 48 higher than that of ViT ($\sim$129-th v.s. $\sim$81-th). Under such premise, compared to ViT, the attention map of iLLaMA can be approximated with a certain error by a higher-rank matrix. Accordingly, the rank of the attention map may affect the expressive capabilities of the learned representations dong2021attention , suggesting that the causal self-attention in iLLaMA has the potential to learn complex visual representations, as empirically demonstrated in Section 4.2. Detailed results for different layers and heads are provided in Appendix D.
@@
Closing remarks. So far, we have finished the design roadmap of iLLaMA through architectural and training strategy modification. iLLaMA, a decoder-only Transformer, shows advantages in computational complexity and attention map rank through its causal self-attention mechanism. Notably, while all components of iLLaMA are essentially derived from LLaMA, relying only on them is insufficient for an effective weight optimization, as demonstrated in Section 3.3. In fact, the proposed PS [cls] and soft mask strategy effectively address this issue and assist in iLLaMA training. However, to achieve a comprehensive understanding of iLLaMA’s properties, some useful evaluation should be conducted: 1) Scalability for large model capacities (>300M parameters) and dataset sizes (>10M training images, e.g., ImageNet-21K). 2) Other practical evaluation dimensions, such as model calibration, shape-texture bias, downstream task performance, quantization compatibility, discussed below.